### v2

# MCP Bridge: A Lightweight, LLM-Agnostic RESTful Proxy for Model Context Protocol Servers

Abstract
Large Language Models (LLMs) are increasingly augmented with external tools through standardized interfaces like the Model Context Protocol (MCP). However, current MCP implementations face critical limitations: they typically require local process execution through STDIO transports, making them impractical for resource-constrained environments like mobile devices, web browsers, and edge computing. We present MCP Bridge, a lightweight RESTful proxy that connects to multiple MCP servers and exposes their capabilities through a unified API. Unlike existing solutions, MCP Bridge is fully LLM-agnostic, supporting any backend regardless of vendor. The system implements a risk-based execution model with three security levels—standard execution, confirmation workflow, and Docker isolation—while maintaining backward compatibility with standard MCP clients. However, reliable execution within this framework requires models that can strictly adhere to protocol schemas. To this end, we also fine-tuned the Qwen3 4B and 8B model family on the Agent-Ark/Toucan-1.5M dataset using four Reinforcement Learning techniques: Group Relative Policy Optimization (GRPO), Dr. GRPO, Beta Normalization Policy Optimization (BNPO), and Decoupled Alignment Policy Optimization (DAPO). Evaluated on the MCPToolBench++ benchmark, our optimized model achieves an F1 score of 73.0% that outperforms GPT-OSS-120B (62.17%) and remains competitive with the 70B+ parameter baselines. Evaluation demonstrates that MCP Bridge successfully addresses the constraints of direct MCP connections while providing enhanced security controls and cross-platform compatibility, enabling sophisticated LLM-powered applications in previously inaccessible environments.
keywords:
Model Context Protocol, Large Language Models, RESTful API, Proxy Architecture, Tool Integration, Risk-Based Execution
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing. They enable sophisticated conversational agents that can understand and generate human-like text across numerous domains [1]. Despite their impressive capabilities, these models are inherently limited by their training data and lack access to real-time information, specialized tools, and the ability to perform actions in external systems [2]. To overcome these limitations, there has been a significant push toward augmenting LLMs with external tools and data sources, allowing them to retrieve information, execute computations, and interact with various services [3].
The Model Context Protocol (MCP) represents a significant advancement in this direction, providing a standardized interface for connecting AI assistants to external tools and data sources [4]. Introduced as an open protocol, MCP aims to establish a universal adapter—a “USB-C port for AI applications”—that enables any compliant model to access any data repository or service through a consistent format. This standardization addresses the fragmentation problem where each new tool integration requires custom development, replacing it with a single, extensible protocol.
However, current MCP implementations face critical limitations that hinder widespread adoption. Many MCP servers rely on STDIO transports that require local process execution, making them impractical for resource-constrained environments such as edge devices, mobile applications, and web browsers. Direct connections to MCP servers from multiple isolated clients also create redundancy and increase resource usage, while the technical complexity of MCP tool formats poses barriers for non-expert users.
In response to these challenges, we present MCP Bridge—a lightweight, fast, and LLM-agnostic proxy that connects to multiple MCP servers and exposes their capabilities through a unified REST API. The architecture is shown in Figure 1. While Anthropic’s MCP SDK provides a reference client/server implementation, MCP Bridge focuses on deployment: it acts as a stable REST adapter that allows heterogeneous clients (e.g., browsers, mobile applications, edge devices) to use MCP servers without local process execution via STDIO transports. MCP Bridge also implements a risk-based execution model that supports standard execution, a confirmation workflow for medium-risk tools, and Docker isolation for high-risk tools, while maintaining backward compatibility with standard MCP clients. The implementation is available as an open-source project at https://github.com/INQUIRELAB/mcp-bridge-api.
Figure 1: Architecture of the MCP Bridge API system showing four layers: client applications (mobile, browser, edge devices, and others) at the top, connecting through a RESTful API to the MCP Bridge proxy, which interfaces with multiple MCP servers (STDIO and SSE) at the bottom. The system enables resource-constrained environments to access MCP functionality through a unified interface with configurable security levels.
Although MCP Bridge is model-agnostic, end-to-end reliability depends on whether the client model can produce MCP-compliant tool calls consistently. To enable open-weight models to act as drop-in MCP Bridge clients, we fine-tune Qwen3-4B and Qwen3-8B using reinforcement learning on Toucan-1.5M [5], comparing GRPO [6], Dr. GRPO [7], DAPO [8], and BNPO [9]. We evaluate the resulting policies on MCPToolBench++ [10] and analyze the role of reward components via ablations (Appendix A).
The remainder of this paper is organized as follows: Section 2 reviews related work in tool-augmented language models and standardization; Section 3 presents the MCP Bridge system design and implementation; Section 4 describes our policy optimization methodology for MCP tool alignment; Section 5 reports experimental results and comparisons; Section 6 discusses future directions; and Section 7 concludes. Reward ablations are provided in Appendix A.
2 Related Work
2.1 Tool Use and Retrieval-Augmented Language Models
Large language models (LLMs) have increasingly been augmented with external data sources and tools to overcome their inherent knowledge and capability limitations [1, 2]. One prominent approach is retrieval-augmented generation (RAG), which integrates a document retriever with the model. Lewis et al. [3] introduced RAG as a general framework combining a parametric neural generator with non-parametric memory of retrieved documents, demonstrating improved performance on knowledge-intensive tasks. By linking to live knowledge sources, retrieval-augmented models can dynamically update their context and provenance, addressing issues like stale knowledge and hallucinations more effectively than static LLMs [3].
Another research direction focuses on enabling LLMs to invoke external tools or APIs. Yao et al.’s ReAct framework [11] interleaves logical reasoning traces with action commands, allowing models to make step-by-step decisions about when to continue thinking or call a tool. Schick et al. extended this concept with Toolformer [12], showing that language models can be self-taught to use tools through training on curated corpora with inserted API calls. HuggingGPT [13] demonstrated model orchestration by using a powerful LLM as a controller that routes user requests to specialized AI models available.
The integration of tools with LLMs has been further advanced through frameworks like Visual ChatGPT [14], which connects LLMs with visual foundation models, and GPT-4Tools [15], which employs self-instruction to teach models new tool-use skills. Lu et al. developed Chameleon [16], a plug-and-play reasoning framework that augments LLMs with modular tools and uses an LLM-based planner to coordinate tool composition. Gorilla [17] addresses real-world API invocation by fine-tuning LLaMA to output exact API calls from a large catalog of machine learning APIs.
2.2 Standardization and LLM-Agnostic Integration
As the ecosystem of LLM-accessible tools expands, integration scalability has emerged as a key challenge—specifically, how to connect any model to any tool with minimal custom code. Anthropic’s Model Context Protocol (MCP) [4] represents a significant standardization effort, introduced as an open standard for connecting AI assistants to software and data. MCP provides a unified client-server architecture where tools and data sources are wrapped as MCP servers that expose specific functions through a consistent interface.
In parallel to protocol standardization, researchers have explored RESTful API proxies as bridges between LLMs and existing web services. RestGPT by Song et al. [18] treats REST APIs as tools that an LLM can learn to use by providing the model with OpenAPI specifications and implementing a coarse-to-fine planning approach. It introduces a dedicated API executor that handles HTTP requests and response parsing, enabling a single framework to access hundreds of different APIs through a uniform method.
Beyond research prototypes, practical frameworks like LangChain and GPT-Index provide unified APIs to access multiple LLM backends and incorporate external tools, reflecting the need for abstraction layers. However, these solutions are largely engineering-focused, whereas academic efforts like MCP aim to establish formal standards that encourage compatibility across diverse systems and vendors.
The MCP Bridge proposed in this work follows this standardization philosophy by acting as a RESTful adapter between MCP servers and client applications. By building on prior ideas of tool augmentation and standard interfaces, MCP Bridge delivers a practical, modular integration solution that remains agnostic to the underlying model or environment. This approach aligns with the broader movement toward making advanced AI functionalities more accessible, interoperable, and future-proof for real-world applications.
3 System Design and Implementation
This section describes the design and implementation of MCP Bridge, a lightweight, fast, and LLM-agnostic proxy for Model Context Protocol (MCP) servers. We detail the system architecture, API design, server management, security model, and client integration components.
3.1 System Architecture and Technology Stack
MCP Bridge follows a layered architecture that decouples client applications from the underlying MCP server processes. Figure 1 illustrates this design, where client applications communicate with the proxy via a standardized REST API, and the proxy manages connections to multiple MCP servers.
The system is built on Node.js (18+) and uses the following core components:
  • Express.js: Provides the HTTP server and routing capabilities
  • Child Process API: Manages spawned MCP server processes
  • Server-Sent Events (SSE): Enables real-time communication between some MCP servers and the proxy
  • Docker SDK: Facilitates containerized execution for high-risk operations
This technology stack was chosen for its minimal footprint, cross-platform compatibility, and non-blocking I/O capabilities—critical requirements for a proxy that must handle multiple concurrent connections with low latency. The implementation uses asynchronous programming patterns throughout to prevent blocking operations from degrading performance.
3.2 RESTful API and Endpoints
MCP Bridge exposes a comprehensive REST API that standardizes access to MCP server functionality. The API is organized into general endpoints for server management and server-specific endpoints for tool execution and resource access.
Table 1 summarizes the primary API endpoints provided by MCP Bridge.
Table 1: MCP Bridge API Endpoints
Endpoint Method Description
/servers GET List all connected MCP servers
/servers POST Start a new MCP server
/servers/{serverId} DELETE Stop and remove a server
/health GET Get health status of MCP Bridge
/confirmations/{id} POST Confirm execution of a medium-risk request
/servers/{id}/tools GET List all tools for a specific server
/servers/{id}/tools/{toolName} POST Execute a specific tool
/servers/{id}/resources GET List all resources
/servers/{id}/prompts GET List all prompts
The API design follows REST principles with JSON as the primary data exchange format. Each endpoint returns appropriate HTTP status codes and standardized error responses. For example, when executing a tool via POST /servers/{id}/tools/{toolName}, the request body contains the tool’s input parameters, and the response includes the execution result or a confirmation request based on the tool’s risk level.
Algorithm 1 API Request Processing Pipeline
1:Input: HTTP request reqreq with server ID sidsid, tool name tooltool, and parameters paramsparams
2:Output: HTTP response resres with result or confirmation request
3:function ProcessToolRequest($req,res$)
4:  $sid\leftarrow req.params.serverId$
5:  $tool\leftarrow req.params.toolName$
6:  $params\leftarrow req.body$
7:  if not ServerExists(sidsid) then
8:   return Error($res,404,\text{"Server not found"}$)
9:  end if
10:  $server\leftarrow\textsc{GetServer}(sid)$
11:  if not ToolExists($server,tool$) then
12:   return Error($res,404,\text{"Tool not found"}$)
13:  end if
14:  $riskLevel\leftarrow\textsc{GetToolRiskLevel}(server,tool)$
15:  if $riskLevel=1$ then
16:   $result\leftarrow\textsc{ExecuteTool}(server,tool,params)$
17:   return Success($res,result$)
18:  else if $riskLevel=2$ then
19:   $confirmationId\leftarrow\textsc{GenerateConfirmationId}$
20:   StoreConfirmationRequest($confirmationId,server,tool,params$)
21:   return RequireConfirmation($res,confirmationId$)
22:  else if $riskLevel=3$ then
23:   $result\leftarrow\textsc{ExecuteToolInDocker}(server,tool,params)$
24:   return Success($res,result$)
25:  end if
26:end function
The request processing pipeline (Algorithm 1) shows how MCP Bridge handles tool execution requests, including validation, risk assessment, and appropriate execution pathways. This unified API layer provides consistent access patterns regardless of the underlying MCP server implementation.
3.3 Server Management and Connection Handling
MCP Bridge dynamically manages connections to MCP servers, supporting both standard STDIO-based servers and newer Server-Sent Events (SSE) implementations. The server management subsystem handles server lifecycle (startup, monitoring, and teardown) and efficiently routes requests to the appropriate server instance.
Algorithm 2 MCP Server Connection Management
1:Input: Server configuration configconfig with command, arguments, and environment variables
2:Output: Server connection object or error
3:function StartMcpServer(configconfig)
4:  $serverId\leftarrow\textsc{GenerateUUID}$
5:  $process\leftarrow\text{null}$ $\triangleright$ Attempt to start MCP server
6:  if $config.command$ exists then
7:   $process\leftarrow\textsc{SpawnProcess}(config.command,config.args,config.env)$
8:  else
9:   return Error: ”Invalid server configuration”
10:  end if
11:  $connected\leftarrow\textsc{WaitForConnection}(process,timeout=5000)$
12:  if not connectedconnected then
13:   KillProcess(process)
14:   return Error: ”Failed to connect to MCP server”
15:  end if
16:  $serverConn\leftarrow\{id:serverId,process:process,tools:[],resources:[]\}$
17:  DiscoverServerCapabilities(serverConn)
18:  RegisterServer(serverId, serverConn)
19:  return serverConnserverConn $\triangleright$ Error handling
20:  if an error occurs then
21:   if $process\neq\text{null}$ then
22:     KillProcess(process)
23:   end if
24:   return Error: error.message
25:  end if
26:end function
As shown in Algorithm 2, server connections are established by spawning child processes or connecting to existing MCP servers via their specified transport. The system automatically discovers each server’s capabilities (available tools, resources, and prompts) upon connection, making them immediately available through the REST API.
The connection manager employs several strategies to maintain robust connections:
  • Heartbeat monitoring: Periodically checks server health
  • Automatic reconnection: Attempts to re-establish lost connections
  • Connection pooling: Optimizes resource usage for high-demand servers
  • Request queueing: Manages concurrent requests to prevent overloading servers
These mechanisms ensure that client applications experience minimal disruption even when underlying MCP servers encounter issues or need to be restarted.
3.4 Security Model and Risk-Based Execution
MCP Bridge implements a comprehensive security model centered around risk-based execution levels. This approach provides granular control over tool invocation, particularly for operations that could potentially modify data or access sensitive resources.
The risk-based execution model defines three levels:
  1. 1.
    Low Risk (Level 1): Standard execution without additional checks, suitable for read-only operations
  2. 2.
    Medium Risk (Level 2): Requires explicit confirmation before execution, appropriate for data-modifying operations
  3. 3.
    High Risk (Level 3): Executed within an isolated Docker container, providing environmental isolation for maximum security
The workflow for medium-risk operations is particularly important (see Algorithm 3), as it implements a two-phase execution pattern that requires explicit confirmation before proceeding.
Algorithm 3 Medium-Risk Confirmation Workflow
1:Input: Confirmation ID confirmationIdconfirmationId, confirmation token tokentoken
2:Output: Execution result or error
3:function ProcessConfirmation($confirmationId,token$)
4:  $pendingReq\leftarrow\textsc{GetPendingRequest}(confirmationId)$
5:  if $pendingReq=\text{null}$ then
6:   return Error: ”Invalid confirmation ID or expired request”
7:  end if
8:  if $pendingReq.token\neq token$ then
9:   return Error: ”Invalid confirmation token”
10:  end if
11:  if $\textsc{IsExpired}(pendingReq)$ then
12:   RemovePendingRequest(confirmationId)
13:   return Error: ”Confirmation expired”
14:  end if
15:  $server\leftarrow\textsc{GetServer}(pendingReq.serverId)$
16:  $result\leftarrow\textsc{ExecuteTool}(server,pendingReq.tool,pendingReq.params)$
17:  RemovePendingRequest(confirmationId)
18:  return resultresult
19:end function
For high-risk operations (Level 3), MCP Bridge leverages Docker containers to provide strong isolation. Each container is configured with specific resource limits, network controls, and volume mounts as defined in the server configuration. This containerization ensures that even if a tool behaves unexpectedly, its impact is contained within the isolated environment.
This multi-tiered approach to security allows system administrators to configure appropriate risk levels based on their security requirements while maintaining compatibility with standard MCP clients that expect direct execution.
3.5 Client Integration (MCP-Gemini Agent)
Complementing the server-side proxy is the MCP-Gemini Agent, a Python client that integrates Google’s Gemini API with MCP Bridge. This agent provides an intelligent natural language interface to the MCP tool ecosystem, allowing users to interact with tools through conversational prompts rather than direct API calls. Figure 2 shows the structure of the LLM integration.
Figure 2: This figure illustrates the integration between a large language model (Gemini) and the Model Context Protocol (MCP) Bridge. The architecture shows a vertical flow from the user through the MCP-Gemini Agent (Python), to the Gemini LLM, then to the MCP Bridge API (Node.js), and finally to the MCP Servers. The diagram highlights the system’s key data flows: natural language inputs from users, structured tool calls from the LLM, and request execution with risk-level handling (low, medium, high). The system supports multi-step reasoning by cycling results back to the LLM to determine subsequent actions.
The MCP-Gemini Agent implements several key features:
  • Multi-step reasoning: Supports complex operations by sequencing tool calls
  • Security confirmation handling: Seamlessly manages the confirmation workflow for medium-risk operations
  • Flexible JSON display: Configurable verbosity for tool outputs
  • Automatic tool discovery: Detects and utilizes all available tools from connected servers
The agent’s architecture follows a conversational loop pattern (see Algorithm 4), where user inputs are processed by the Gemini LLM to generate appropriate tool calls to MCP Bridge.
Algorithm 4 MCP-Gemini Agent Conversation Loop
1:Input: User query queryquery, MCP Bridge URL mcpUrlmcpUrl
2:Output: Agent response with tool execution results
3:function ProcessUserQuery($query,mcpUrl$)
4:  $toolList\leftarrow\textsc{FetchAvailableTools}(mcpUrl)$
5:  $prompt\leftarrow\textsc{BuildToolAwarePrompt}(query,toolList)$
6:  $llmResponse\leftarrow\textsc{InvokeGeminiLLM}(prompt)$
7:  $tools\leftarrow\textsc{ExtractToolCalls}(llmResponse)$
8:  $results\leftarrow$ empty list
9:  for each tooltool in toolstools do
10:   $result\leftarrow\textsc{ExecuteMcpTool}(mcpUrl,tool.serverId,tool.name,tool.params)$
11:   if $result.requiresConfirmation$ then
12:     $confirmation\leftarrow\textsc{PromptUserForConfirmation}(result.confirmationDetails)$
13:     if $confirmation.confirmed$ then
14:      $result\leftarrow\textsc{ConfirmToolExecution}(mcpUrl,result.confirmationId)$
15:     else
16:      $result\leftarrow\{status:"cancelled",message:"Usercancelledoperation$"\}
17:     end if
18:   end if
19:   Append resultresult to resultsresults
20:  end for
21:  $followupPrompt\leftarrow\textsc{BuildResultPrompt}(query,tools,results)$
22:  $finalResponse\leftarrow\textsc{InvokeGeminiLLM}(followupPrompt)$
23:  return finalResponsefinalResponse
24:end function
The MCP-Gemini Agent demonstrates that proprietary LLMs can use MCP Bridge without additional training by emitting structured tool calls directly; in Section 4, we show that open-weight models benefit from explicit RL alignment to satisfy MCP’s strict formatting and tool-selection requirements.
The agent implementation is modular and configurable, supporting various command-line options for customizing behavior:
usage: llm_test.py [-h] [--hide-json] [--json-width JSON_WIDTH]
                   [--mcp-url MCP_URL] [--mcp-port MCP_PORT]
MCP-Gemini Agent with configurable settings
options:
  -h, --help            show this help message and exit
  --hide-json           Hide JSON results from tool executions
  --json-width JSON_WIDTH
                        Maximum width for JSON output (default: 100)
  --mcp-url MCP_URL     MCP Bridge URL (default: http://localhost:3000)
  --mcp-port MCP_PORT   Override port in MCP Bridge URL
Beyond demonstrating MCP Bridge with a proprietary LLM client, the proxy is designed to be model-agnostic: any language model capable of emitting MCP-style structured tool calls can use the same REST endpoints to discover tools, invoke them, and receive results. In the next section, we describe how we fine-tune open-weight Qwen3 models to produce MCP-compliant tool calls reliably, enabling them to act as drop-in clients of MCP Bridge.
4 Methodology: Policy Optimization for Tool Alignment
This section describes how we align open-weight Qwen3 models for reliable MCP-style tool use via reinforcement learning, while keeping the MCP Bridge interface model-agnostic. We focus on two practical objectives: selecting the correct tool(s) and emitting a parseable MCP-compliant tool-call structure that MCP Bridge can execute.
4.1 Problem Setup and Output Interface
Each example consists of a natural-language query $q$, a set of available tools $\mathcal{T}$ (tool name and description, optionally schema), and a reference set of expected tool names $\mathcal{T}^{\star}$. The model produces a response $y$ that may include one or more tool calls. In training and evaluation, tool calls are expected to appear as JSON enclosed in explicit tags:
<tool_call>
{"name": "tool_name", "arguments": {"arg1": "...", "arg2": "..."}}
</tool_call>
We extract the predicted tool set $\widehat{\mathcal{T}}(y)$ from the model output and compare it to $\mathcal{T}^{\star}$. Tool calls are extracted using robust pattern matching over (i) tagged <tool_call> blocks and (ii) common JSON-like formats produced by LLMs during tool use. To avoid false mismatches arising from server prefixes and naming conventions, tool names are normalized by lowercasing, removing known server prefixes (e.g., Server::tool_name $\rightarrow$ tool_name), stripping common MCP prefixes/suffixes, and canonicalizing separators (hyphens/underscores/spaces) before matching.
4.2 Training and Evaluation Data
We train on tool-use data derived from Toucan-1.5M [5]. The training script filters the SFT portion to subsets that contain tool calls (specifically single-turn-original, single-turn-diversify, and multi-turn) [5]. Each training instance provides a user question, a tool list tools (with per-tool name and description), and a comma-separated target_tools field indicating the expected tool(s). During prompt construction, the tool list is inserted into the system message under an Available Tools header, with the number of tools and description length capped to reduce context length pressure.
We evaluate on MCPToolBench++ [10], a large-scale benchmark organized by tool category (Browser, File System, Search, Map, Pay, Finance). Each evaluation sample provides a query, tool metadata, and ground-truth function-call labels specifying the expected tool name(s). For evaluation, each category is scored independently, and we report Precision, Recall, F1, and Accuracy aggregated across samples.
4.3 Metrics
Let $P=\widehat{\mathcal{T}}(y)$ be the predicted tool set after normalization and $G=\mathcal{T}^{\star}$ be the ground-truth tool set. We compute Precision $|P\cap G|/|P|$, Recall $|P\cap G|/|G|$, and $F1=2\cdot\text{Prec}\cdot\text{Rec}/(\text{Prec}+\text{Rec})$ (with standard zero-handling). Accuracy is defined as exact set match (i.e., $P=G$) for single-turn MCPToolBench++ tasks.
4.4 Reward Signal Used for RL
Policy optimization uses a scalar reward that combines (i) whether the model chose the correct tool(s) and (ii) whether the tool call is formatted in a parseable MCP-compatible structure. Concretely, the total reward is:
$r(y;q,\mathcal{T})\;=\;r_{\text{sel}}(y)\;+\;r_{\text{fmt}}(y).$
The tool-selection component $r_{\text{sel}}$ is derived from tool-set overlap (Precision/Recall/F1) between the extracted tool set and the reference tool set; the training script uses a piecewise mapping that strongly rewards near-perfect F1 and penalizes missing or completely incorrect tool calls. The format component $r_{\text{fmt}}$ rewards generating a <tool_call> block containing valid JSON with the required name field, and assigns lower or negative scores when the output is malformed or not parseable.
The training code also includes additional reward variants (e.g., stricter JSON validity checks, schema compliance, parameter completeness, and length penalties), but our experiments found that the best-performing configuration for the models reported in this paper uses only the tool-selection reward and the format reward.
4.5 Policy Optimization Methods
We compare four policy-optimization variants implemented in the training pipeline: GRPO [6], Dr. GRPO [7], DAPO [8], and BNPO [9]. All methods operate on on-policy samples and update the model using group-based relative advantages (multiple generations per prompt), which reduces variance without requiring an explicit learned critic. GRPO [6] provides the baseline group-relative objective. Dr. GRPO [7] modifies the objective to address training pathologies observed in R1-zero-like recipes. DAPO [8] introduces an asymmetric clipping variant (with an additional high clipping parameter) designed to preserve learning signal as policies improve. BNPO [9] introduces a bounded normalization strategy that stabilizes updates under shifting reward distributions.
4.6 Implementation Details (Training Configuration)
We implement training using Unsloth [19] for efficient Qwen3 loading and LoRA adaptation, and TRL-style policy optimization with group sampling. Table 2 summarizes the configuration used in our training script.
Table 2: Training configuration used for Qwen3 tool-use policy optimization in our pipeline.
Setting Value
Backbone models Qwen3-4B and Qwen3-8B (Unsloth checkpoints) [19]
Max sequence length 2048
Quantization 4-bit loading enabled
LoRA rank ($r$) 32
LoRA target modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj
LoRA alpha $2r$ (64)
Optimizer / scheduler AdamW 8-bit, linear schedule with warmup ratio 0.1
Learning rate / weight decay $5\times 10^{-6}$ / 0.001
Batching per-device batch size 1, gradient accumulation 1
On-policy generations 4 samples per prompt (group size)
Sampling parameters temperature 1.0; top-$p$ 1.0; min-$p$ 0.1; EOS stop enabled
Training steps / checkpointing max steps 500; save every 200 steps
Reward functions used tool-selection + format rewards (other reward variants tested but not used in final runs)
5 Experiments and Results
We evaluate tool-use behavior on MCPToolBench++ [10], reporting Precision, Recall, F1, and Accuracy for six tool categories (Browser, File System, Search, Map, Pay, Finance). We compare the base Qwen3 models against four RL methods (GRPO, Dr. GRPO, DAPO, BNPO), trained on tool-use data derived from Toucan-1.5M [5] using the tool-selection and format rewards described in Section 4.
5.1 Per-category F1 trends
Figure 3 summarizes per-category F1 scores. Two patterns appear consistently across model sizes. First, File System is the strongest category for all methods. Second, Map is persistently challenging across methods, with F1 clustered in a narrow band; this indicates that simply improving selection/formatting does not fully address category difficulty, likely due to higher semantic ambiguity in map-related intent and overlapping tool affordances.
Figure 3: Per-category F1 score heatmaps on MCPToolBench++ for Qwen3-4B (top) and Qwen3-8B (bottom). Rows correspond to RL methods and the base model; columns correspond to tool categories. Each cell reports mean F1 (%) computed from normalized tool-name matching and robust tool-call extraction from model outputs. Darker green indicates better performance.
A radar view of the same F1 values is shown in Figure 4. For Qwen3-4B, Dr. GRPO improves overall balance across Browser, Search, and Pay relative to the base model, whereas DAPO shows mixed gains with a notable drop in Finance. For Qwen3-8B, BNPO expands the performance most strongly in Browser while maintaining competitive Pay and Search, while Dr. GRPO produces the highest Finance score.
Figure 4: Radar plots of per-category F1 on MCPToolBench++. This view emphasizes cross-category trade-offs: improvements are not uniform across tools, and methods can help one category while degrading another (e.g., Finance sensitivity in smaller models).
5.2 Aggregate metrics and method comparison
Figure 5 reports aggregate Precision, Recall, F1, and Accuracy (averaged over categories). For Qwen3-4B, Dr. GRPO yields the strongest aggregate F1 and Accuracy, improving tool selection reliability relative to the base model while preserving high File System performance. GRPO is close behind, indicating that group-relative updates alone deliver meaningful gains when the reward signal is dominated by selection correctness and format validity. In contrast, DAPO and BNPO do not consistently outperform Dr. GRPO at 4B, suggesting that small models are more sensitive to reward distribution shifts and clipping choices, particularly in Finance.
For Qwen3-8B, BNPO achieves the highest aggregate Precision/Recall/F1 and the best Accuracy, indicating that as capacity increases, bounded/normalized updates can convert the same reward signal into more consistent improvements across categories. Notably, Finance improves substantially for 8B under GRPO-family methods, implying that higher-capacity policies can better internalize tool-call patterns and disambiguate tool choice even when category semantics are difficult. An ablation study isolating the tool-selection reward and the format reward is provided in Appendix A.
Figure 5: Aggregate Precision, Recall, F1, and Accuracy (%) on MCPToolBench++ [10] for Qwen3-4B (top) and Qwen3-8B (bottom). Scores are averaged across the six tool categories. These results show that Dr. GRPO is the strongest overall method for Qwen3-4B, while BNPO is strongest for Qwen3-8B under the same reward design.
5.3 Comparison against other models
Table 3 compares our best MCP-aligned checkpoints to several representative baselines using the same metric suite. The comparison contextualizes the tool-use reliability of MCP-aligned Qwen3 models against larger models that may have broader general capabilities. While the largest baselines achieve higher absolute scores, the RL-aligned Qwen3 models narrow the gap substantially relative to their base versions, demonstrating that targeted policy optimization with selection+format rewards can produce practical tool-use behavior at smaller scales.
Table 3: Tool-use evaluation comparison (Precision/Recall/F1/Accuracy).
Model Precision (%) Recall (%) F1 (%) Accuracy (%)
MCP Bridge (Qwen3-4B + Dr. GRPO) 66.0 66.7 66.2 65.3
MCP Bridge (Qwen3-8B + BNPO) 72.2 74.7 73.0 69.7
GPT-OSS-120B 61.11 64.33 62.17 58.00
Kimi-K2-Instruct 76.83 80.00 77.89 73.67
Llama-3.3-70B 81.50 82.00 81.67 81.00
Llama-4-Maverick-17BA-400B 78.18 80.67 78.94 76.00
Llama-4-Scout-17BA-109B 76.36 77.67 76.67 75.67
6 Future Work
MCP Bridge establishes a practical foundation for LLM-agnostic tool integration, and our RL-aligned Qwen3 clients demonstrate that open-weight models can reliably operate within this interface. Several directions remain to strengthen both the deployment layer and the model-alignment layer.
On the systems side, MCP Bridge can be extended with production-oriented features such as stronger multi-tenancy support, richer observability (structured tracing of tool calls and outcomes), and adaptive caching for idempotent requests. While the current implementation supports STDIO and SSE-backed MCP servers, broader transport coverage and improved server lifecycle management (including persistent server pools and backpressure-aware scheduling) would improve throughput under heavy load. Security can be expanded beyond the current risk levels by introducing fine-grained access policies (e.g., per-tool or per-argument allow/deny rules), integrating standard authentication/authorization mechanisms for enterprise deployments, and supporting more specialized sandbox profiles for high-risk tool execution.
On the modeling side, our current RL signal focuses on correct tool selection and MCP-compliant formatting. A natural next step is to incorporate execution-grounded feedback where rewards reflect tool-call success, argument validity, and downstream task completion rather than tool identity alone. This includes learning to fill parameters accurately, handling tool errors robustly, and improving multi-step tool planning over longer horizons. Additionally, aligning policies to the risk-based execution model itself (e.g., learning when a user confirmation is required or how to request missing permissions) could further improve safety and usability when operating against real MCP server fleets.
Finally, we plan to broaden evaluation beyond tool-name matching to include end-to-end success metrics, multi-turn tool-use trajectories, and real-world MCP environments. This will help quantify generalization to unseen servers, unseen tool schemas, and compositional tool sequences, which are critical for deploying MCP-based agents at scale.
7 Conclusion
This paper introduced MCP Bridge, a lightweight, LLM-agnostic RESTful proxy that addresses key limitations of direct connections to Model Context Protocol (MCP) servers—especially the reliance on local STDIO execution that is impractical for browsers, mobile clients, and edge deployments. MCP Bridge exposes MCP server capabilities through a unified API, supports both STDIO and SSE server backends, and provides a risk-based execution model with standard execution, confirmation workflows, and Docker isolation for high-risk tools.
Beyond the proxy layer, we showed that reliable MCP usage depends on the client model’s ability to produce tool calls that are both correct and protocol-compliant. To support open-weight deployment, we fine-tuned Qwen3-4B and Qwen3-8B using four RL methods (GRPO, Dr. GRPO, DAPO, BNPO) on Toucan-1.5M and evaluated them on MCPToolBench++. Across model sizes, RL improves tool-use reliability relative to the base models, with Qwen3-4B + Dr. GRPO achieving 66.2% F1 and 65.3% Accuracy, and Qwen3-8B + BNPO achieving 73.0% F1 and 69.7% Accuracy. These results outperform the GPT-OSS-120B baseline in our comparison table and narrow the gap to much larger 70B+ parameter baselines.
Taken together, MCP Bridge and RL-aligned open-weight clients provide a practical path toward deployable MCP-based agents: the proxy makes MCP server fleets accessible across platforms, and the aligned policies make open models reliable protocol clients. We hope this combination helps move MCP from a local-development interface toward a broadly deployable tool-access layer for heterogeneous AI systems.
Declarations
Funding: This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Clinical Trial Number: Not applicable.
Consent to Publish: Not applicable.
Data Availability: The complete MCP Bridge implementation is available as an open-source project at https://github.com/INQUIRELAB/mcp-bridge-api.
Ethics and Consent to Participate: Not applicable.
Competing Interests: The authors declare that they have no competing interests.
Appendix A Reward Ablation Study
This appendix isolates the effect of the two reward components used in our RL training: the tool-selection reward and the format reward. We compare two ablations—Tool Select. Only and Format Only—against the main setting used in Section 5 (tool-selection + format). Across both Qwen3-4B and Qwen3-8B, the ablations show that the two reward components are complementary: optimizing formatting alone can improve protocol compliance and yield moderate metric gains, while optimizing selection alone provides limited benefit and can be unstable for smaller models. The best overall performance is achieved when both rewards are used jointly in the main experiments.
Table 4: Summary of best aggregate results (%) under reward ablations (macro-averaged over categories).
Model Full (Sel.+Fmt.) Tool Select. Only Format Only
Qwen3-4B (best F1 / Acc.) 66.2 / 65.3 63.1 / 61.3 65.8 / 64.0
Qwen3-8B (best F1 / Acc.) 73.0 / 69.7 71.6 / 68.7 72.3 / 68.3
Figure 6: Per-category F1 heatmaps under the Tool Select. Only ablation for Qwen3-4B (top) and Qwen3-8B (bottom). This ablation evaluates how much performance can be recovered when the reward focuses only on selecting the correct tool(s), without explicitly rewarding MCP-compliant formatting.
Figure 7: Radar plots of per-category F1 under the Tool Select. Only ablation. The plots highlight that selection-only training does not uniformly improve categories and can degrade specific domains (notably Finance for smaller models), motivating the joint reward used in the main experiments.
Figure 8: Aggregate Precision/Recall/F1/Accuracy under the Tool Select. Only ablation. Compared to the main setting (Section 5), selection-only training yields modest gains at 8B and limited or negative gains at 4B, indicating that explicit formatting incentives are important for stable improvements.
Figure 9: Per-category F1 heatmaps under the Format Only ablation for Qwen3-4B (top) and Qwen3-8B (bottom). This ablation rewards emitting a parseable MCP tool-call structure without directly optimizing tool identity, testing how much of tool-use reliability comes from enforcing output structure alone.
Figure 10: Radar plots of per-category F1 under the Format Only ablation. Format-only optimization improves consistency in several categories, but the strongest results still come from jointly optimizing both selection and formatting (Section 5).
Figure 11: Aggregate Precision/Recall/F1/Accuracy under the Format Only ablation. Format-only training can produce sizable gains (especially for Qwen3-4B), but combining format and tool-selection rewards yields the best overall metrics and the most reliable tool-use behavior.